# Chaos in hyperscaling violating Lifshitz theories

### v1

@@ IV   Entanglement wedge
Consider some perturbation in the boundary theory. As time passes, the information from the perturbation gets delocalized over a larger and larger region, i.e. the size of the applied operator increases. The size of the operator is the smallest region that contains the information of the applied perturbation operator. In the dual bulk picture, it corresponds to the extremal surface ”just” enclosing the particle (i.e. the particle is at the tip of the extremal surface at all times) (refer figure 2). It’s important to note that the background geometry is static and it’s only the position of the particle that is time-dependent. So we use the RT surface [44] as our extremal surface and not the HRT surface [45]. For our case, the entanglement wedge is a constant t hyper-surface, bounded by the RT surface and the boundary sub-region.
@@
Here $\gamma$ is the determinant of the induced metric on the co-dimension 2 surface in the bulk and $\xi$ are the co-ordinates on the surface. Extremising the above action (or equivalently the induced metric) will give rise to the equation of the RT surface. So we first compute the induced metric on the surface. Since the surface is embedded in a $t=constant$ hypersurface, the tt component of the induced metric vanishes. Further, we take $\tilde{r}$ to be the radial co-ordinate in the $x^{i}$ direction ($\tilde{r}=|x|$) and parametrize r with $\tilde{r}$ ($r=r(\tilde{r})$). Hence, from (1) we get the following induced metric.
@@
In order to compute the butterfly velocity, we do a near-horizon analysis of the RT surface (refer figure 2). Near the horizon, we can take the following form of r($\tilde{r}$)
@@
For simplicity, here we have taken the horizon radius $r_{h}=1$, $\epsilon$ is some arbitrarily small parameter and u($\tilde{r}$) is the function that we need to determine by extremizing (44). Substituting (46) in the induced metric and expanding near the horizon, upto $\mathcal{O}(\epsilon)$, we get the following.
@@
[…]
Figure 2: The growth of the entanglement wedge (shaded in yellow) as the particle, originating from the boundary (represented by the red dot), propagates towards the horizon. The figure on the left shows the entanglement wedge at some time $t$ and the figure on the right shows the entanglement wedge at a sufficiently later time $t^{\prime}$, when the RT surface reaches near the horizon and the near-horizon profile of the RT surface is given by $u(r)$.
Here $f_{o}$ represents the blackening function f(r) evaluated at the horizon $r_{h}=1$. The prime denotes the derivative with respect to $\tilde{r}$. Extremising (47) (and thereby $S_{EE}$) we arrive at the following differential equation in u($\tilde{r}$).
@@
$u(r)\sim\frac{e^{\mu\tilde{r}}}{\tilde{r}^{n}}$ (49)
@@
The above result matches the result obtained via the shockwave analysis (42). Note that for $\theta=0$, we recover the result obtained in [38].
V   Variation of $v_{B}$ with z, $\theta$, Q and $r_{h}$
$v_{B}$ for HVL theories (given by (39)) exhibits a rich dependence on various parameters (particularly $\theta$ and z). In this section, we point out several interesting features (that can provide insights into the chaotic behavior of the HVL theories), which are by no means exhaustive. It is important to note that the constraints on $\theta$ and z that arise from the null energy condition are crucial for this analysis. To reiterate, for planar horizon topology, there is no black hole solution for $z<1$, for solution to exist for $21\leq z<$ 2, $\theta\leq(d-1)(z-1)$ and for $z\geq 2$, $\theta<d-1$. Since Lorentz invariance is broken in the HVL theories, we will observe that $v_{B}$ is not bounded, which is why all the plots in this section correspond to $v_{B}<1$. For the analysis in this section, we fix $d=4$ and since $v_{B}$ is symmetric in Q, we only study $Q>0$. Also note that, we discuss the permissible combination of the four parameters {z,$\theta$, Q, $r_{h}$}. By permissible we mean that the combination of the parameters yield a $v_{B}$ that is real, a temperature that is positive and the constraints on $\theta$ and z are satisfied.
V.1   Variation of $v_{B}$ with z
$v_{B}$ exhibits several interesting features, when varied with respect to z, keeping $\theta$, Q and $r_{h}$ fixed. In this subsection, we fix $\theta=1$ and study the behavior of $v_{B}$ with respect to z for different values of $r_{h}$ and Q.
(a) Here, $r_{h}=1$ and $\theta=1$. The red, blue and orange lines correspond to $Q=0.5$ (<1)<1)< 1 ), $Q=1$ and $Q=1.3(>1)$ respectively.
(b) Here $r_{h}=0.9(<1)$ and $\theta=1$. The red and the orange curves correspond to $Q=0.1$ and $Q=0.6$ respectively. The blue dot represents the maxima that occurs at $z=2.643$
(c) Here $r_{h}=1.1(>1)$ and $\theta=1$. The red and the orange curves correspond to $Q=1.8$ and $Q=1.3$ respectively. The blue dot represents the minima that occurs at $z=2.124$
@@ V.1   Variation of $v_{B}$ with z
Since $\theta=1$ and $d=4$, for $21\leq z<$ 2, we have that $z\geq 4/3$ and for $z\geq 2$, the constraint is satisfied automatically by the choice of $\theta$. This implies that the allowed values of z are $[\frac{4}{3},\infty)$. We study three sub-cases, namely $r_{h}=1$, $r_{h}<1$ and $r_{h}>1$. Within each sub-case we study how the charge Q affects $v_{B}^{2}(z)$.
For $r_{h}=1$, $v_{B}^{2}$ is a linear function of z. For $Q=1$, $v_{B}^{2}$ becomes constant, whereas for $Q>1$, $v_{B}^{2}$ decreases with increase in z and for $Q<1$, $v_{B}^{2}$ increases in the allowed range of z values.
For $r_{h}\neq 1$, $v_{B}^{2}$ varies non-linearly with z. For both $r_{h}<1$ and $r_{h}>1$, there are certain (not all) permissible combinations of {$r_{h}$, Q} that lead to a non-monotonic variation of $v_{B}^{2}$ with respect to z (in the allowed range) and certain combinations that yield a monotonic behavior. Lets consider an example for the subcase $r_{h}<1$. For $r_{h}=0.9$ and $Q=0.1$, we get a maxima of $v_{B}^{2}$ at $z=2.643$. Whereas for $r_{h}=0.9$ and $Q=0.6$, we get a monotonically decreasing behavior. Now consider the subcase $r_{h}>1$. For $r_{h}=1.1$ and $Q=1.8$, we get a minima of $v_{B}^{2}$ at $z=2.124$, whereas for same $r_{h}=1.1$ but different charge $Q=1.3$, $v_{B}^{2}$ monotonically increases with z. The difference between the two subcases is that, for $r_{h}<1$, we get a maxima in the non-monotonic case and a decreasing behavior in the monotonic case, whereas for $r_{h}>1$, we get a minima and an increasing behavior respectively. (One can check that this statement is true for all permissible combinations).
The combinations of parameters discussed in this subsection (and in the following subsections) indeed yield a positive temperature. (one can check that by plotting T on the same plot as $v_{B}^{2}$.)
V.2   Variation of $v_{B}$ with $\theta$
Just like z, variation of $v_{B}$ with respect to $\theta$ also yields some interesting results. Here, we consider $21\leq z<$ 2 and $z\geq 2$ separately because the two set of z values are accompanied by different constraints. From the first set (i.e. $21\leq z<$ 2), we pick $z=1.8$, which gets accompanied by the constraint $\theta\leq 2.4$. From the second set ($z\geq 2$) we
(a) Here $r_{h}=1$ and $z=1.8$. The red, orange and blue curves correspond to $Q=0.8$, $Q=1$ and $Q=1.2$ respectively.
(b) Here $r_{h}=0.5(<1)$, $z=1.8$. The red, orange, blue and green curves correspond to $Q=0$, $Q=0.01$, $Q=0.1$ and $Q=0.3$ respectively.
(c) Here $r_{h}=1.1(>1)$ and $z=1.8$. The red, orange and blue curves correspond to $Q=1$, $Q=1.4$ and $Q=1.8$ respectively. The green dots represent the minimas. Corresponding to $Q=1.4$ and $Q=1.8$, the minimas occur at $\theta=-0.59$ and $\theta=0.305$. Note that we have suppressed the respective maximas that occur at $\theta=-21.43$ and $\theta=-25.8$ respectively.
@@ V.2   Variation of $v_{B}$ with $\theta$
can pick any value of z (because the corresponding constraint does not depend on z) and the constraint is $\theta<3$ (since $d=4)$. We again deal with the three sub-cases ($r_{h}=1$, $r_{h}<1$ and $r_{h}>1$) and analyze how the various values of Q affect the variation of $v_{B}$ with respect to $\theta$.
For $r_{h}=1$, $v_{B}^{2}$ monotonically increases with $\theta$ for all permissible values of Q and z. Note that the monotonically increasing behavior also holds for $\theta<0$.
For $r_{h}<1$ also, $v_{B}^{2}$ increases with increasing $\theta$. It is worth noting that for $Q=0$, $v_{B}$ is real and $T>0$ for all the allowed values of $\theta$. (for both $z\in[1,2)$ and $z\geq 2$). However, just a slight increase in the charge, significantly cuts down the range of $\theta$ values for which $v_{B}$ is real and $T>0$. For example, consider $z=1.8$ and $r_{h}=0.5$. For $Q=0$, $\theta\in(-\infty,2.4]$, for $Q=0.01$, $\theta\in[-3,2.4]$ and for $Q=0.1$, $\theta\in[0.08,2.4]$ (this is shown in figure 4(b)).
For $r_{h}>1$, $v_{B}^{2}$ exhibits both monotonic and non-monotonic behaviors. It is interesting to note that we get two extrema in this case. We get a minima which can occur at both $\theta<0$ or $\theta>0$ and we get a maxima which occurs at $\theta<0$. Whether we get a monotonic or non-monotonic behavior, depends on the corresponding values of $r_{h}(>1)$, z and Q. For both $z\in[1,2)$ and $z\geq 2$, there exist values of $r_{h}(>1)$ and Q for which non-monotonic behavior is obtained. For example, for $z=1.8$ $r_{h}=1.1$ and $Q=1.8$, we get a minima at $\theta=0.305$ and a maxima at $\theta=-25.803$. Whereas for same z and $r_{h}$ but different charge $Q=1.4$, we get a minima at $\theta=-0.591$ and a maxima at $\theta=-21.43$.
V.3   Variation of $v_{B}$ with Q
The variation of $v_{B}^{2}$ with Q is much simpler. For any permissible value of $r_{h}$ of z and $\theta$, $v_{B}^{2}$ monotonically decreases with Q (i.e. $Q>0$. We essentially mean that $v_{B}^{2}$ decreases with an increase in the magnitude of the charge).
@@ V.3   Variation of $v_{B}$ with Q
(a) Here, $r_{h}=1$ and $z=1.5$. The red, orange and blue curves correspond to $\theta=1.2$, $\theta=0.7$ and $\theta=0.2$ respectively.
@@
(b) Here, $r_{h}=1$ and $\theta=1$. The red, orange and blue curves correspond to $z=1.4$, $z=1.6$ and $z=1.8$ respectively.
Figure 5: $v_{B}^{2}$ vs Q
V.4   Variation of $v_{B}$ with $r_{h}$
The variation of $v_{B}^{2}$ with $r_{h}$ is also simple. It monotonically increases with increase in $r_{h}$ for all permissible values of Q values of z and $\theta$.
@@ V.4   Variation of $v_{B}$ with $r_{h}$
(a) Here $Q=1.5$ and $z=1.5$. The red, orange and blue curves correspond to $\theta=1.2$, $\theta=0.7$ and $\theta=0.2$ respectively.
@@
(b) Here $Q=1.5$ and $\theta=1$. The red, blue and orange curves correspond to $z=1.4$, $z=1.6$ and $z=1.8$ respectively.
Figure 6: $v_{B}^{2}$ vs $r_{h}$
VI   Discussion and future outlook
@@
In this work we presented a detailed discussion on the shockwave analysis in HVL theories and thereby computed the various chaos parameters. $v_{B}$ was also computed using the entanglement wedge method. Note that, our analysis was restricted to planar horizon topology. One can also do a similar analysis for hyperbolic and spherical topologies and analyze what affect does k ($\neq 0$) has, on the behavior of $v_{B}$.
We pointed out various interesting features in the behavior of $v_{B}$ when varied with respect to z, $\theta$, Q and $r_{h}$. When varied with z (keeping $\theta$, $r_{h}$, and Q fixed) and $\theta$ (keeping z, $r_{h}$, and Q fixed), non-monotonicities are obtained for certain values of the corresponding fixed parameters. Unpacking these non-monotonicities can perhaps provide deeper insights into the chaotic nature and the dynamics of the HVL theories.
Moreover, $v_{B}$ was found to monotonically decrease with increasing Q for all permissible choices of z, $\theta$ and $r_{h}$. This behavior of $v_{B}$ with respect to Q is also obtained in one of author’s upcoming works which deals with completely different class of holographic theories. Thus, this behavior of $v_{B}$ might be a universal feature of all holographic theories.
Also, $v_{B}$ is found to increase with an increase in the horizon radius ($r_{h}$). For HVL geometries, the entropy S $\sim r_{h}^{d-\theta-1}$ and since $\theta<d-1$, S increases with increase in $r_{h}$. This implies that entropy increases the rate of scrambling, irrespective of the values of z, $\theta$ and Q. One might suspect that this is also a universal feature of all holographic theories.
A key direction for future work is to develop a concrete understanding of the findings presented here.
VII   Acknowledgements

### v2

@@ IV   Entanglement wedge
Consider some perturbation in the boundary theory. As time passes, the information from the perturbation gets delocalized over a larger and larger region, i.e. the size of the applied operator increases. The size of the operator is the smallest region that contains the information of the applied perturbation operator. In the dual bulk picture, it corresponds to the extremal surface ”just” enclosing the particle (i.e. the particle is at the tip of the extremal surface at all times) (refer figure 2). It’s important to note that the background geometry is static and it’s only the position of the particle that is time-dependent. So the RT surface [46] is used as the extremal surface and not the HRT surface [47]. Thus, in this case the entanglement wedge is a constant t hyper-surface, bounded by the RT surface and the boundary sub-region.
@@
Here $\gamma$ is the determinant of the induced metric on the co-dimension 2 surface in the bulk and $\xi$ are the co-ordinates on the surface. Extremising the above action (or equivalently the induced metric) will give rise to the equation of the RT surface. So we first compute the induced metric on the surface. Since the surface is embedded in a $t=constant$ hypersurface, the tt component of the induced metric vanishes. Further, taking $\tilde{r}$ to be the radial co-ordinate in the $x^{i}$ direction ($\tilde{r}=|x|$) and parametrizing r with $\tilde{r}$ ($r=r(\tilde{r})$), we get the following induced metric from (1)
@@
In order to compute the butterfly velocity, a near-horizon analysis of the RT surface is done (refer figure 2). Near the horizon, the following form of r($\tilde{r}$) can be taken.
@@
[…]
Figure 2: The growth of the entanglement wedge (shaded in yellow) as the particle, originating from the boundary (represented by the red dot), propagates towards the horizon. The figure on the left shows the entanglement wedge at some time $t$ and the figure on the right shows the entanglement wedge at a sufficiently later time $t^{\prime}$, when the RT surface reaches near the horizon and the near-horizon profile of the RT surface is given by $u(\tilde{r})$.
For simplicity, here we have taken the horizon radius $r_{h}=1$, $\epsilon$ is some arbitrarily small parameter and u($\tilde{r}$) is the function to be determined by extremizing (47). Substituting (49) in the induced metric and expanding near the horizon, upto $\mathcal{O}(\epsilon)$, we get the following.
@@
Here $f_{o}^{\prime}$ represents the derivative of the blackening function, evaluated at the horizon. Extremising (50) (and thereby $S_{EE}$) we arrive at the following differential equation in u($\tilde{r}$).
@@
$u(\tilde{r})\sim\frac{e^{\mu\tilde{r}}}{\tilde{r}^{n}}$ (52)
@@
The above result matches the result obtained via the shockwave analysis (45).
V   $v_{B}$ in terms of the boundary thermodynamic variables
The next goal is to understand how boundary thermodynamics and theory parameters ($\theta$ and z) affect information scrambling in HVL theories. In order to achieve that goal, we first express $v_{B}$ (given by (42)) purely in terms of boundary thermodynamic variables, using a recently developed thermodynamic dictionary for HVL theories [38]. This dictionary has various merits, such as the matching of the Smarr formula on the bulk side with the Euler relation on the boundary side, as well as the matching of the first laws on the boundary and in the bulk. Here, we list the dictionary entries that are relavant for expressing $v_{B}$ in terms of the boundary thermodynamics. The entries for central charge, entropy, and charge are as follows
$C=\frac{\Omega_{k,d-1}L^{d-\theta-1}r_{F}^{\theta}}{16\pi G_{N}}$ (57)
$S=4\pi Cx^{d-\theta-1}$ (58)
$Q=\sqrt{2Z_{o}}Cq\hskip 2.84544pt[(d-\theta-1)(d-\theta+z-3)]^{\frac{1}{2}}$ (59)
Here, x is a bulk parameter given as $x\equiv\frac{r_{h}}{L}$ and S and Q are dimensionless. Note that, in none of the relavant entries (for our purposes), is there an appearance of the arbitrary length parameter $r_{o}$ (refer [38]) and thus there is no effect of $r_{o}$ on how information scrambles in the boundary. We continue working with $L=1$ (and $k=0$). $r_{h}$ and q are the two bulk parameters that we need to eliminate in (42) and thus we express these parameters in terms of the boundary quantities in the following way.
$r_{h}=\tilde{S}^{\frac{1}{d-\theta-1}}$ (60)
$q=\frac{\tilde{Q}}{[(d-\theta-1)(d-\theta+z-3)]^{1/2}}$ (61)
where $\tilde{S}$ and $\tilde{Q}$ are defined as
$\tilde{S}\equiv\frac{S}{4\pi C}$ (62)
$\tilde{Q}\equiv\frac{Q}{\sqrt{2Z_{o}}C}$ (63)
Substituting (60) and (61) in (42), we get the following expression for $v_{B}$ in terms of the boundary thermodynamic variables.
$v_{B}^{2}=\frac{\tilde{S}^{\frac{2(z-1)}{d-\theta-1}}(d-1+z-\theta)(d-\theta-1)-\left(\tilde{Q}/\tilde{S}\right)^{2}}{2(d-\theta-1)^{2}}$ (64)
We treat $Z_{o}$ as a constant parameter. It is interesting to note that the entropy couples to both z and $\theta$, however charge does not couple to z. With the expression (64) in hand, we now study the variation of $v_{B}$ with respect to z, $\theta$, $\tilde{S}$ and $\tilde{Q}$.
VI   Variation of $v_{B}$ with z, $\theta$, $\tilde{S}$ and $\tilde{Q}$
In this section, we study the variation of $v_{B}$ and note some interesting features. Before proceeding, there are certain important points to highlight. Given that the Lorentz invariance is broken in HVL theories, $v_{B}$ is not bounded by $v_{B}^{S}$ [45], where $v_{B}^{S}$ is the butterfly velocity for AdS-Schwarzschild black brane, given as
$v_{B}^{S}=\sqrt{\frac{d}{2(d-1)}}$ (65)
and is, moreover, super-luminal for certain values of the parameters. Secondly, in this section, we only deal with the ”permissible” values of the parameters and thermodynamic variables, where by permissible we mean the combination {z, $\theta$, $\tilde{S}$, $\tilde{Q}$ for which the temperature $T>0$. Furthermore, we also keep track of the constraints on z and $\theta$ imposed by the null energy condition, discussed in II.1. For analysis in this section, we fix $d=4$.
First, to analyze only the effect of z, we set $\theta=0$ (i.e. purely Lifshitz theories). Initially, we deal with the case where $\tilde{Q}=0$ and then consider the effect of non-zero charge. For $\theta=0$ and $\tilde{Q}=0$ ,
(a) $\tilde{S}<1$, $\tilde{Q}=0$, $\theta=0$, and $d=4$ for this plot. The blue, red and green curves correspond to $\tilde{S}=0.7$, $\tilde{S}=0.75$ and $\tilde{S}=0.8$ respectively. The black dots represent the non-monotonicities.
(b) $\tilde{S}\geq 1$, $\tilde{Q}=0$, $\theta=0$ , and $d=4$ for this plot. The red, blue and green curves correspond to $\tilde{S}=1$, $\tilde{S}=1.5$ and $\tilde{S}=2$ respectively.
@@ VI   Variation of $v_{B}$ with z, $\theta$, $\tilde{S}$ and $\tilde{Q}$
(c) $\tilde{Q}=0$, $z=1$, and $d=4$ for this plot. The plot is the same for all values of $\tilde{S}$.
(d) $\tilde{Q}\neq 0$, $z=1$, $\tilde{S}=1$ and $d=4$ for this plot. The red, blue, and green curves correspond to $\tilde{Q}=1.4$, $\tilde{Q}=1.5$ and $\tilde{Q}=1.6$. The black dots represent the non-monotonicities.
@@
(a) $\tilde{Q}=1$, $\theta=1$, and $d=4$ for this plot. The green, blue and the red curves correspond to $z=3$, $z=2.5$, and $z=2$ respectively.
@@
(b) $\tilde{Q}=1$, $z=2$, and $d=4$ for this plot. The green, blue and the red curves correspond to $\theta=1.5$, $\theta=1.2$, and $\theta=1$ respectively.
Figure 5: $v_{B}^{2}$ vs $\tilde{S}$
@@
(a) $\tilde{S}=1.1$, $\theta=0.5$, and $d=4$ for this plot. The red, blue and green curves correspond to $z=1.2$, $z=1.3$, and $z=1.4$ respectively.
@@
(b) $\tilde{S}=1.1$, $z=1.5$, and $d=4$ for this plot. The red, blue and green curves correspond to $\theta=0.6$, $\theta=0.8$, and $\theta=1$ respectively.
Figure 6: $v_{B}^{2}$ vs $\tilde{Q}$
$v_{B}$ interestingly exhibits non-monotonic behavior with respect to z when $10<\tilde{S}<$ 1, in the permissible region. We get a maxima. However, only in the range $0.8770.688<\tilde{S}<$ 0.877, the non-monotonicities lie in the allowed region for z, i.e. ($z>1$) and $v_{B}$ at the non-monotonic point is not superluminal. Since the charge does not couple to z, there is no non-trivial effect of non-zero $\tilde{Q}$. $\tilde{Q}\neq 0$ merely shifts the values and does not change the behavior of $v_{B}$ with respect to z. For $\tilde{S}\geq 1$, $v_{B}$ monotonically increases with z in the permissible region. This suggests that $v_{B}$ does not necessarily increase with z in Lifshitz theories, as reported in earlier works (such as [39]), but it depends on entropy per central charge.
Next, to study the effect of $\theta$, we set $z=1$ (i.e. purely hyperscaling-violating theories). For $z=1$, the corresponding constraint on $\theta$ is $\theta\leq 0$. We divide the discussion into two subcases $\tilde{Q}=0$ and $\tilde{Q}\neq 0$. For $z=1$ and $\tilde{Q}=0$, $v_{B}$ monotonically increases with $\theta$ for all values of $\tilde{S}$. However, since charge couples with $\theta$, the effect of non-zero charge is non-trivial. $\tilde{Q}\neq 0$ induces a non-monotonicity in $v_{B}$ when varied with respect to $\theta.$ We get a maxima (again, we mean in the permissible region, for allowed values of $\theta$ with $v_{B}$ not being superluminal).
It is important to note that the non-monotonicities in $v_{B}$, as a function of $z$ and $\theta$, do not coincide with those in temperature. This can be verified by plotting $T$ and $v_{B}$ on the same graph. Hence, the underlying causes of the non-monotonicities in $v_{B}$ and temperature are not the same.
With respect to $\tilde{S}$, $v_{B}$ increases monotonically in the permissible region and for all allowed values of $z$ and $\theta$. Note that the entropy dependence of $v_{B}$ arises due to the appearance of $r_{h}$ in (36). However, $v_{B}$ is a model-dependent parameter, and in certain holographic models, $r_{h}$ does not appear in the expression for $v_{B}$. In such models, entropy does not affect information scrambling. One example of this is the planar RNAdS black hole.
On the other hand, $v_{B}$ monotonically decreases with respect to $\tilde{Q}$ (by which we mean the magnitude of the charge) in the permissible region and for all allowed values of z and $\theta$. It is interesting to note that, in one of the author’s upcoming works, quantum chaos in a completely different class of holographic models has been studied, and it is found that $v_{B}$ necessarily depends on charge and moreover, in all the models studied in the work, $v_{B}$ is found to monotonically decrease with charge (or more precisely, the magnitude of the charge). This observation suggests that the slowing down of the rate of information scrambling by charge might be a universal feature of all holographic models or even a characteristic of all quantum systems in general.
VII   Discussion and future outlook
@@
In this work, we presented a detailed discussion on the shockwave analysis in HVL theories and computed various chaos parameters. $v_{B}$ was also computed using the entanglement wedge method, and the results obtained from both methods were found to match. Note that our analysis was restricted to planar horizon topology. One can extend the analysis to hyperbolic and spherical topologies and analyze what effect a non-zero $k$ has on the behavior of $v_{B}$. However, this task would require devising certain new methods.
Furthermore, using a recently developed thermodynamic dictionary, we expressed $v_{B}$ purely in terms of boundary thermodynamics, which yielded a neat result. We then discussed the variations of $v_{B}$ with respect to $z$, $\theta$, $\tilde{S}$, and $\tilde{Q}$. Notably, we observed a non-monotonic behavior of $v_{B}$ when varied with respect to $z$ for $\tilde{S}<1$, whereas $v_{B}$ was found to increase monotonically with $z$ for $\tilde{S}\geq 1$. This implies that $z$ does not necessarily increase $v_{B}$ (as reported in previous works [39]). This suggests that entropy non-trivially affects the rate of information scrambling. Thus, to gain deeper insights into the nature of information scrambling in any holographic model, it is more useful to retain $r_{h}$ in the expression of $v_{B}$ (instead of setting it equal to one, as is commonly practiced in the literature).
In uncharged hyperscaling-violating theories, $v_{B}$ was found to increase monotonically with $\theta$, whereas for non-zero charge, a non-monotonic behavior was obtained. Additionally, $v_{B}$ was observed to increase monotonically with $\tilde{S}$ and decrease monotonically with $\tilde{Q}$. The latter fact also holds for a completely different class of holographic models discussed in one of the author’s upcoming works, and thus one can conjecture that charge decreases the rate of information scrambling in all holographic models or quantum systems in general. It would be interesting to investigate whether a counter-example to this conjecture exists.
When varying $z$ or $\theta$, one navigates through a space of theories. The observed non-monotonicities suggest that the theories on either side of the non-monotonic points may possess distinguishing features, leading to different natures of information scrambling on either side. Unpacking the underlying physics behind these non-monotonicities (as well as other features discussed) is an interesting direction for future work. Additionally, we noted that $v_{B}$ equals the rate of entanglement growth during thermalization, indicating a connection between the two, which would also be an interesting avenue to explore further.
VIII   Acknowledgements